OpenStudy (hamzajani):

Equation of normal to the parabola y = 5 - x 2 at the point where x = 2 is

sam (.sam.):

What do you mean "2 at the point"?

OpenStudy (hamzajani):

Y=5-x^2

sam (.sam.):

I don't understand your question, is it the equation of normal to the parabola is y = 5 - x^2, at the point where x = 2 is?

OpenStudy (hamzajani):

Yes

OpenStudy (canimcan):

Jednacina od normale do parabole?

jhonyy9 (jhonyy9):

@Hamzajani are you hungarian ?- sorry that i ask you in this way but i not can sending to nobody messages bc. the submit button of my profile was/is blocked

OpenStudy (mathmale):

y = 5 - x 2 at the point where x = 2 is Please format your equation properly. If the given equation is a parabola, then it must have the form y = c + bx + ax^2, or y - ax^2 + bx + c. First find the slope of the tangent line to the graph of this function at x=2. Then the slope of the line normal to the graph at x = 2 is the negative reciprocal of the slope of the tangent line. Please show the work you have done and ask questions, if need be, so that someone can help you finish finding the equation of the normal line.